Consider the following reaction: Cr2O72– + 6Br– + 14 H+ 2 Cr3+ + 3Br2 + 7H2O
In a redox titration, 14.62 mL of 0.175 M Cr2O72– was needed to completely react with a 25.00 mL sample of Br–. Calculate the [Br–] in the original solution.
1
Expert's answer
2020-06-02T14:07:32-0400

According to the reaction equation for the interaction of 1 equivalent of dichromate, we will need 6 equivalents of bromide ions. According to the condition of the question we have 0.175*14.62 = 2.56 mmol of dichromate. For full interaction of this quantity we need 2.56*6 = 15.35 mmol of Br. Respectively, [Br] in the original solution is = 15.35/25.00 = 0.614 mol/L.

Answer: [Br] = 0.614 M